Explain to grade 4 class why pencils on the table are 3/10 of the pencils, not 3/7


1
Expert's answer
2021-06-07T13:05:31-0400

Complete question:We have three pencils on the table, other seven pencils are on the floor. Explain to a grade 4 class why pencils on the table are 3/10 of the pencils, not 3/7


When we have 3 pencils on the table and rest of the 7 pencils are on the floor.


Then the total number of pencils "=" pencils on table "+" pencils on the floor

total number of pencils "=3+7=10"


Now Factor of pencils on the table = pencils on the table divided by total pencils

According to this:

Factor of pencils on the table "= \dfrac{3}{10}"( of the total pencils)


The answer can not be "\dfrac{3}{7}" .as denominator always contains the total no of pencils present(that is 10 and not 7).